10m^2-23m+12=0 It says to find the solution set.
Answer by flame8855(424)

10m^2-23m+12=0
a=10 b=-23 c =12
delta= (b^2-4ac)^2= (23^2-40(12))=49
x=(23+7)/20=1.5
x=(23-7)/20=0.8